# Geometry and cohomology of compactified Deligne–Lusztig varieties

### v1

@@ 2.2. The Lang map
There is an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
Definition 2.5.
Let $F_{1}:G^{r}\to G^{r}$ be the partial Frobenius morphism given by
@@
Similarly, we define the partial Frobenius morphism on $\left(G/B\right)^{r}$. Since $B$ is the base change of $B_{0}$ to ${\rm Spec}\ k$, we know that $G/B$ is the base change of $G_{0}/B_{0}$ to ${\rm Spec}\ k$ and has a corresponding $q$-Frobenius map $F:G/B\to G/B$. Denote by $f_{2}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map given by
$(x_{0},\ldots,x_{r-2},x_{r-1})\longmapsto(x_{0},\ldots,x_{r-2},F(x_{r-1})),$
which is the identity on the first $(r-1)$-entries and Frobenius on the last entry. Denote by $f_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ the map
$(x_{0},\ldots,x_{r-1})\longmapsto(x_{1},\ldots,x_{r-1},x_{0}),$
which shifts the coordinates by $1$. The composition $f_{1}^{r}$ is the identity.
Definition 2.6.
We define the partial Frobenius morphism $F_{1}:\left(G/B\right)^{r}\to\left(G/B\right)^{r}$ to be the composition $f_{2}\circ f_{1}$.
The partial Frobenius morphism $F_{1}$ on $\left(G/B\right)^{r}$ is an étale morphism and the composition $F_{1}^{r}$ is the $q$-Frobenius morphism on $\left(G/B\right)^{r}$.
Lemma 2.7.
The Lang map $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$.
Proof.
Note that $G$ is geometrically connected and smooth, so the product $G^{r}$ is connected and smooth. Observe that the composition of the partial Frobenius $F_{1}^{r}$ yields the Frobenius on $G^{r}$, so [CE, Theorem 7.1] applied to $G^{r}$ yields that $\mathcal{L}_{1}$ is surjective and étale. The kernel of $\mathcal{L}_{1}$ is $\Delta(G^{F})$, where $\Delta:G\to G^{r}$ is the diagonal embedding. Thus $\mathcal{L}_{1}$ is a finite morphism.
Since the diagonal (left) action of $G^{F}$ is a free action and induces automorphisms of the cover, and the number of automorphisms is the same as the degree of $\mathcal{L}_{1}$, we know that $\mathcal{L}_{1}$ is a Galois covering with Galois group $G^{F}$ [SP, 03SF]. ∎
Lemma 2.8.
@@
Furthermore, $G^{F}$ acts transitively on the connected components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equi-dimensional.
@@
We record the proof for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ as the proof for $X(w_{1},\ldots,w_{r})$ is the same. In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via the projection ${\rm pr}_{0,\ldots,r-1}:\left(G/B\right)^{r+1}\to\left(G/B\right)^{r}$ to the first $r$-entries. The diagonal right $({B})^{r}$-action on $G^{r}$ is free. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the right $({B})^{r}$-action, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. The isomorphism above then follows from verifying the definitions cf. [DL]. With the diagonal $G^{F}$-action on $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$, the isomorphism is $G^{F}$-equivariant.
We show the second assertion for $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Via Lemma 2.7 we know that
@@
is irreducible, so $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of connected components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Therefore $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equidimensional. ∎
Remark 2.9.
@@ 2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BR], Görtz [Gör], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BR].
@@
Definition 2.10.
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{k}(w)|k\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
@@
An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$. For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
@@
Proposition 2.12.
[…]
@@
Take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$. This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
@@
via projection. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove connectedness for ${\overline{X}^{S^{*}}(v)}$.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BR] for classical Deligne–Lusztig varieties. Suppose ${{\text{supp}}}_{F}(w_{1},\ldots,w_{r})=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ has the same expression as $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [Lus], [BR].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible.
@@
2.4. Criterion for smoothness
In [Wan, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties.
Proposition 2.13.
For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F$-rational and thus normal and Cohen-Macaulay.
@@ 2.4. Criterion for smoothness
The proof is similar to the proof of [Wan, Proposition A.6]. For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F$-regular [BT, Cor. 4.1, p. 958]. Thus the product
@@
is strongly $F$-regular by faithfully flat descent [HH, Theorem 5.5]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F$-regular cf. [Smi, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Abe, Theorem 3.6] and conclude that
@@
is strongly $F$-regular. Finally, by Lemma 2.8 and the faithfully flat descent of strong $F$-regularity [HH, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F$-regular. Strongly $F$-regular rings are $F$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HH, Theorem 4.2, 6.27]. ∎
Remark 2.14.
@@
Since $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal by 2.13, the connectedness criterion for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ in 2.12 applies readily for irreducibility.
Since smoothness can be verified étale locally and satisfies faithfully flat descent, the same proof as in 2.13 gives an alternative proof for the smoothness result in Lemma 2.3.
Corollary 2.15.
Let $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ be as in 2.13. If for all $i=1,\ldots,r$, the Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth. In particular, ${\overline{X}^{S^{*}}(w)}$ is smooth for all $w\in S^{*}$.
Example 2.16.
If for all $i=1,\ldots,r$, $w_{i}$ is the longest element in $W_{{\text{supp}}_{F}(w_{i})}$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@ 3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups cf. [GP].
All the relations in the group presentation of the Weyl group (as a finite Coxeter group $(W,S)$) can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (3.1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (3.1) a braid relation in the Weyl group $W$. The classification of finite groups of Lie type ([DM, Classification 4.3.6]) yields that for $s\neq t$, we have $m_{st}\in\{2,3,4,6\}$.
Remark 3.1.
Whenever $(st)^{m_{st}}=1,\ s,t\in S$, is a braid relation in $W$,
$(ts)^{m_{st}}=1$ (3.2)
also holds. This is a consequence of the relation $ss=1$ for all $s\in S$. We also have relations
$(F(s)F(t))^{m_{st}}=1\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$ (3.3)
corresponding to the braid relation (3.1). When there is no ambiguity, we refer to any equation among (3.1), (3.2) and (3.3) as a braid relation in $W$.
Definition 3.2.
Let $w\in S^{*}$. If $w=w_{1}w_{2}$, for some $w_{1},w_{2}\in S^{*}$, we say that $w_{i}$ divides $w$, $i=1,2$. Moreover, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.3.
@@
$P_{m}(s,t):=st\cdots\in S^{*}$ (3.4)
such that ${{\ell^{S^{*}}}(P_{m}(s,t))}=m$ and $P_{m}(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P_{m}(s,t)$ to denote its canonical projection $\alpha(P_{m}(s,t))\in W$.
Lemma 3.4.
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $i\in\{1,\ldots,r\}$, we have $w_{i}=P_{m}(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@
By Lemma 2.3 (resp. Corollary 2.15), it suffices to show that for all $i=1,\ldots,r$, the associated Schubert variety $\overline{{B}w_{i}{B}}/{B}$ is smooth.
We show this by using the classification of finite groups of Lie type and the classification of finite Coxeter groups. Denote $I=\{s,t\}\subseteq S$, $W_{I}\subseteq W$ the corresponding subgroup generated by $I$, and $P_{I}\subseteq G$, $L_{I}\subseteq P_{I}$ the parabolic and Levi subgroups respectively. Note that $L_{I}$ is reductive and $P_{I}/{B}$ is isomorphic to $L_{I}/({B}\cap L_{I})$, which is a flag variety and thus is smooth. We proceed by showing that $S(P_{m}(s,t))=P_{I}/{B}$.
Since $W_{I}$ is generated by two elements, the longest element is determined by the braid relation. When $L_{I}$ is of type $A_{1}\times A_{1}$ (resp. ${{\ell^{S^{*}}}(P_{2}(s,t))}=2$), $st\in W_{I}$ is the longest element. When $L_{I}$ is of type $A_{2}$ or ${}^{2}A_{2}$ (resp. ${{\ell^{S^{*}}}(P_{3}(s,t))}=3$), $sts\in W_{I}$ is the longest element. When $L_{I}$ is of type $B_{2}$ or ${}^{2}B_{2}$ (resp. ${{\ell^{S^{*}}}(P_{4}(s,t))}=4$), $stst\in W_{I}$ is the longest element. When $L_{I}$ is of type $G_{2}$ or ${}^{2}G_{2}$ (resp. ${{\ell^{S^{*}}}(P_{6}(s,t))}=6$), $ststst\in W_{I}$ is the longest element. Therefore $S(P_{m}(s,t))=P_{I}/{B}$ for $m=2,3,4,6$. ∎
Definition 3.5.
Denote by $B^{+}$ the monoid generated by $S$ and all the relations of the form
$P_{m}(s,t)=P_{m}(t,s)$
with $s,t\in S$, $t\neq s$, such that $(st)^{m}=1$ is a braid relation in $W$.
More details about this monoid can be found in [GP, §4].

### v2

@@ 2.2. The Lang map
We consider an alternative description of Deligne–Lusztig varieties via the Lang map. Using this description, we show that $G^{F}$ acts transitively on the set of connected components of $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$. Denote by $G^{r}$ the $r$-fold product $G\times\cdots\times G$ over ${\rm Spec}\ k$. This is again a $k$-group scheme.
We begin with fixing notations. Let $F_{1}:G^{r}\to G^{r}$ be the endomorphism given by
@@
Note that $F_{1}$ can be viewed as a Frobenius root for the algebraic group $G^{r}$.
@@
Lemma 2.2.1.
@@
Furthermore, $G^{F}$ acts transitively on the irreducible components of $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$). In particular, $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are equidimensional.
@@
We record the proof for $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ as the proof for $X(w_{1},\ldots,w_{r})$ is the same. In this proof, we denote $\mathcal{B}:=\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}$.
By Lang’s theorem as formulated in [CabanEngue, Theorem 7.1], we deduce that $\mathcal{L}_{1}:G^{r}\to G^{r}$ is surjective and is a Galois cover with Galois group $G^{F}$. The subscheme $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ of $G^{r}$ is stable under the free $({B})^{r}$-action from the right, so we may take the quotient of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)$ with respect to the induced free right $({B})^{r}$-action. It follows that $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is a closed subscheme of $\left(G/B\right)^{r}$. Observe that $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is also isomorphic to a closed subscheme of $\left(G/B\right)^{r}$ via projection to the first $r$-entries. The isomorphisms in this lemma then follows from verifying the definitions, cf. [DL, §1.11].
Moreover,
@@
is irreducible. Since the quotient map $\pi:\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)\to\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is equivariant under the left $G^{F}$-action and surjective, it induces a surjective morphism $G^{F}\backslash(\mathcal{L}_{1}^{-1}(\mathcal{B}))\to G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$. Thus $G^{F}\backslash\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$ is irreducible and $G^{F}$ acts transitively on the set of irreducible components of $\mathcal{L}_{1}^{-1}\left(\mathcal{B}\right)/({B})^{r}$.
Remark 2.2.2.
@@ 2.3. Criterion for connectedness
There is a simple criterion for classical Deligne–Lusztig varieties to be connected, proved by Lusztig, Bonnafé–Rouquier [BonnafeRouquierOntheirred], Görtz [GoertzOntheconn], and many others. In this section we prove that $X(w_{1},\ldots,w_{r})$ (resp. $X(\underline{w_{1}},\ldots,\underline{w_{r}})$) is connected when the tuple $(w_{1},\ldots,w_{r})$ has full $F$-support by modifying the proof of [BonnafeRouquierOntheirred].
@@
Definition 2.3.1.
The $F$-orbit of $w\in W$ is the set ${{\rm Orb}_{F}(w)}:=\{F^{n}(w)|n\in\mathbb{Z}_{\geq 0}\}$. Under the assumption that $F(S)=S$, we define the set
@@
An element $w\in W$ is called of full $F$-support if ${{\text{supp}}}_{F}(w)=S_{F}$.
For a tuple $(w_{1},\ldots w_{r})\in W^{r}$, we denote
@@
Proposition 2.3.3.
Let $w_{1},\ldots,w_{r}\in W$ and denote $I={{\text{supp}}}_{F}(w_{1},\ldots,w_{r})$. The Deligne–Lusztig varieties $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are irreducible if and only if $I=S_{F}$. In general, they have $|G^{F}/P_{I}^{F}|$ connected components, where $W_{I}$ is the $F$-stable parabolic subgroup of $W$ generated by $\{s\in W|{{\rm Orb}_{F}(s)}\in I\}$ and $P_{I}={B}W_{I}{B}$.
@@
For certain reduced expressions of $w_{1},\ldots,w_{r}$, take the product $w_{1}\cdots w_{r}$ in $S^{*}$ and denote it by $v$. This product depends on the choices of the reduced expressions of $w_{1},\ldots,w_{r}$, but for each such $v$, there is an isomorphism
@@
via projection, cf. [DMR, §2.3.1]. It also follows that $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are dense open subschemes of ${\overline{X}^{S^{*}}(v)}$. Thus it suffices to prove that ${\overline{X}^{S^{*}}(v)}$ is connected if and only if $v$ has full $F$-support.
The proof showing that ${\overline{X}^{S^{*}}(v)}$ is connected is analogous to the proof of [BonnafeRouquierOntheirred] for classical Deligne–Lusztig varieties. Suppose $I=S_{F}$ and thus ${{\text{supp}}}_{F}(v)=S_{F}$. Let $Z$ be the connected component of ${\overline{X}^{S^{*}}(v)}$ which contains the point $({B},\ldots,{B})$. Denote the stabiliser of $Z$ in $G^{F}$ by ${\rm Stab}(Z)$. Fix $v^{\prime}$ to be of minimal length such that $v^{\prime}\preceq v$ and ${{\text{supp}}}_{F}(v^{\prime})=S_{F}$. In particular, $v^{\prime}$ is a reduced expression of $\alpha(v^{\prime})\in W$, which is a Coxeter element. This implies that ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible [LusztigCoxeterOrbits, Proposition 4.8].
On the other hand, there is a $G^{F}$-equivariant embedding ${\overline{X}^{S^{*}}(v^{\prime})}\to{\overline{X}^{S^{*}}(v)}$. Let $Z^{\prime}$ be the connected component of ${\overline{X}^{S^{*}}(v^{\prime})}$ containing $({B},\ldots,{B})$. Then ${\rm Stab}(Z^{\prime})\subseteq{\rm Stab}(Z)$. However, ${\overline{X}^{S^{*}}(v^{\prime})}$ is irreducible, so ${\rm Stab}(Z^{\prime})=G^{F}$. Thus ${\overline{X}^{S^{*}}(v)}$ is irreducible. In this case, its dense open subschemes $X(w_{1},\ldots,w_{r})$ and $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ are also irreducible [GoerWedhAlgGeom1, Proposition 3.24].
@@
2.4. $F_{p}$-regularity
We denote by $F_{p}$ the (absolute) $p$-Frobenius to distinguish our notations from the notations coming from the theory of $F_{p}$-singularities in commutative algebra and algebraic geometry.
In [Wang22, Proposition A.6], we showed that some compactified Deligne–Lusztig varieties are strongly $F_{p}$-regular. This property generalises to arbitrary compactified Deligne–Lusztig varieties. As in loc. cit., we use a similar strategy as in the proof of [Lusztigunipotentch, Lemma 4.3], which uses faithfully flat descent of the smoothness property along the Lang map.
Proposition 2.4.1.
For $w_{1},\ldots,w_{r}\in W$, the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular. In particular, $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is $F_{p}$-rational and thus normal and Cohen-Macaulay.
If $\overline{Bw_{i}B}$ is smooth for all $i$, then $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@ 2.4. $F_{p}$-regularity
When $\overline{Bw_{i}B}$ is smooth for all $i$, this is proved by faithfully flat descent of smoothness in [Lusztigunipotentch, Lemma 4.3] and [DMR, Proposition 2.3.5], cf. [GoerWedhAlgGeom1, Proposition 14.59]. Analogously, we use faithfully flat descent of strongly $F_{p}$-regularity to show that any compactified Deligne–Lusztig variety is strongly $F_{p}$-regular.
For each $i=1,\ldots,r$, we know that $\overline{{B}w_{i}{B}}$ is strongly $F_{p}$-regular [BrTh, Cor. 4.1, p. 958] and thus Cohen-Macaulay [BrTh, Cor. 4.2, p. 959]. The product
@@
is strongly $F_{p}$-regular by faithfully flat descent along the projection map ${\rm pr}_{i}:\overline{{B}w_{1}{B}}\times_{{\rm Spec}\ k}\cdots\times_{{\rm Spec}\ k}\overline{{B}w_{r}{B}}\to\overline{{B}w_{i}{B}}$ [HoHu, Theorem 5.5.b]. Recall that the Lang map $\mathcal{L}_{1}$ is étale, so its closed fibres are smooth and thus strongly $F_{p}$-regular, cf. [KSmith, Ch 3, Remark 1.6, Theorem 4.4(a)]. We may now apply [Aber, Theorem 3.6] and conclude that
@@
is strongly $F_{p}$-regular. Finally, by Lemma 2.2.1 and the faithfully flat descent of strong $F_{p}$-regularity [HoHu, Theorem 5.5], $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is strongly $F_{p}$-regular.
Strongly $F_{p}$-regular rings are $F_{p}$-rational, so $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is normal and Cohen-Macaulay by [HoHu, Theorem 4.2, 6.27]. ∎
Remark 2.4.2.
@@ 3.1. Smoothness via braid relations
We consider some specific elements in the Weyl group for which the compactified Deligne–Lusztig variety is smooth. To begin, we recall the relations in the Weyl group of finite reductive groups, cf. [GeckPfeiffer].
All the relations in the group presentation of the Weyl group as a finite Coxeter group $(W,S)$ can be written as
$(st)^{m_{st}}=1,\ s,t\in S,$ (1)
where $m_{st}\in\mathbb{Z}$ such that $m_{ss}=1$ and $m_{st}\geq 2$ if $s\neq t$. In the case of $s\neq t$, we call (1) a braid relation in the Weyl group $W$. We refer to such an equation (1) as a braid relation in $W$.
Remark 3.1.1.
If $F(S)=S$, then whenever $(st)^{m_{st}}=1,\ s,t\in S$ is a braid relation in $W$, the equations
$(ts)^{m_{st}}=1,\hskip 14.22636pt\ (F(s)F(t))^{m_{st}}=1,\hskip 14.22636pt\text{and}\hskip 14.22636pt(F(t)F(s))^{m_{st}}=1$
also hold.
For $w=w_{1}w_{2}\in S^{*}$, we call $w_{1}$ a prefix of $w$ and $w_{2}$ a suffix of $w$.
Definition 3.1.2.
@@
$P(s,t):=st\cdots\in S^{*},$ (2)
such that ${{\ell^{S^{*}}}(P(s,t))}=m$ and $P(s,t)$ is a prefix of $(st)^{m}$ in $S^{*}$. When there is no ambiguity, we also use $P(s,t)$ to denote its canonical projection $\alpha(P(s,t))\in W$.
Lemma 3.1.3.
Let $w_{1},\ldots,w_{r}\in W$. Suppose that for all $j\in\{1,\ldots,r\}$, we have $w_{j}=P(s,t)$ for some braid relation in $W$. Then the compactified Deligne–Lusztig variety $X(\underline{w_{1}},\ldots,\underline{w_{r}})$ is smooth.
@@
By Proposition 2.4.1, it suffices to show that for all $j=1,\ldots,r$, the associated shceme $\overline{{B}w_{j}{B}}$ is smooth.
Consider $w_{j}=P(s,t)$ with $(st)^{m}=1$. Denote $J=\{s,t\}\subseteq S$. The the corresponding subgroup $W_{J}\subseteq W$ generated by $J$ has a group presentation
$W_{J}=\langle s,t|(st)^{m}=1,s^{2}=1,t^{2}=1\rangle.$
As a consequence, $P(s,t)$ is a reduced expression of the longest element of $W_{J}$.
Let $P_{J}=BW_{J}B\subseteq G$ be the parabolic subgroup corresponding to $W_{J}$. Since $w_{j}$ is the longest element in $W_{J}$, we know that $\overline{{B}w_{j}{B}}=P_{J}$ and is smooth cf. [DigneMichelBook, Proposition 3.2.8]. ∎